THE ν_{1} AND ν_{6} BANDS OF DIIODOMETHANE, CH_{2}I_{2}, AROUND 3.3 MICRONS STUDIED BY HIGH-RESOLUTION FOURIER-TRANSFORM SPECTROSCOPY

J. ORPHAL, N. IBRAHIM, Laboratoire Interuniversitaire des Systémes Atmosphériques (LISA), CNRS UMR 7583, Université de Paris-XII, 91405 Créteil Cedex, France; C. E. FELLOWS, Laboratório de Espectroscopia e Laser, Instituto de Física, Universidade Federal Fluminense, Niterói, Rio de Janeiro, Brazil.

Diiodomethane, CH_{2}I_{2}, is an important molecule in marine photochemistry. CH_{2}I_{2} is produced by algae^{1} and its photolysis is a source for tropospheric iodine which is very reactive towards other organic species and ozone; it is also an important source for marine particle formation^{2}. CH_{2}I_{2} is a rather heavy molecule and has low-lying vibrational modes, leading to a very dense spectrum. Its ground^{3} and first excited^{4} vibrational states have been studied in the past using Fouriertransform microwave spectroscopy. Very recently, quantitative infrared absorption spectra of CH_{2}I_{2} were recorded at a spectral resolution of 0.1 cm^{−1} as part of the PNNL database of gas-phase infrared spectra.^{5} In this paper we present the first high-resolution spectra of the ν_{1} and ν_{6} bands of CH_{2}I_{2} in the 3.3 μm region, recorded with a Bruker IFS-120 HR Fourier-transform spectrometer, using CH_{2}I_{2} at its saturated vapour pressure at room temperature, with an absorption path of 300 cm and a spectral resolution of 0.002 cm^{−1}.

[1] C. O’Dowd et al., Nature, 417, 632-636 (2002).

[2] J. L. Jimenez et al., J. Geophys. Res. D, 108, 4090, doi:10.1029/2002JD002452 (2003).

[3] Z. Kisiel et al., J. Chem. Phys., 105, 1778-1785 (1996).

[4] Z. Kisiel et al., J. Mol. Spectrosc., 199, 5-12 (2000).

[5] T. J. Johnson et al., Atmos. Chem. Phys. Discuss., 6, 1275-1299 (2006).